OpenStudy (anonymous):

Xlogy-logz.ylogz-logx.zlogx-logy=1 how to prove this??

OpenStudy (anonymous):

it was supposed to x to the power logy-logz y to the power logz-logx z to the power logx-logy

hartnn (hartnn):

\(\huge x^{\log y-\log z}y^{\log z-\log x}z^{\log x-\log y}\) like this ?

OpenStudy (anonymous):

ah thank god..yes like that

hartnn (hartnn):

consider \(\huge \log[x^{\log y-\log z}y^{\log z-\log x}z^{\log x-\log y}]\) can u simplify that ?

hartnn (hartnn):

use, \(\log AB=\log A+\log B \\ \log A^{B}=B\log A\)

hartnn (hartnn):

you should get that =0 = ln 1

OpenStudy (anonymous):

i dont understand y u multiplied the whole by log????

hartnn (hartnn):

to be able to simplify

hartnn (hartnn):

else, nothing can't be really done to that expression

OpenStudy (anonymous):

i have done that problem in this way \[x ^{\log y/z}y ^{\log z/x}z ^{\log x/y}\]

OpenStudy (anonymous):

and i cudnt figure out the next step..... :(

hartnn (hartnn):

i know, i tried that also....since i could not proceed, i thought of taking log....

OpenStudy (anonymous):

if u dont mind,can u show me the simplification process...??? :)

hartnn (hartnn):

\(\huge \log[x^{\log y-\log z}y^{\log z-\log x}z^{\log x-\log y}] \\\large =\log[x^{\log y-\log z}]+\log[y^{\log z-\log x}]+\log[z^{\log x-\log y}]\) got this ?

hartnn (hartnn):

now use \( \\ \log A^{B}=B\log A\)

hartnn (hartnn):

@suss

OpenStudy (anonymous):

oh sorry man i jus lost connection leeme check it out

OpenStudy (anonymous):

y did u keep the + sign ???

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

because \(\log AB=\log A+\log B \\ \)

OpenStudy (anonymous):

ah finally got it .....thx a ton man

hartnn (hartnn):

sure ? u getting that = 0 ? then write 0=log 1

OpenStudy (anonymous):

yes 100%..i hav to use formula of log ab bak ther ......i really got it man

hartnn (hartnn):

good :) welcome ^_^